# Node Perturbation Can Effectively Train Multi-Layer Neural Networks

### v6

@@ 2.2 Increasing perturbation learning efficiency through decorrelation
Decorrelation can be combined with any of the formulations described in the previous sections by replacing $\mathbf{x}_{l}$ in Eq. 1 with $\mathbf{x}_{l}^{*}$. We will use DBP, DNP, DINP and DANP when referring to the decorrelated versions of the described learning rules.
In Appendix E, Algorithm 1, we describe decorrelation using the decorrelated activity-based update procedure.
2.3 Experimental validation
To measure the performance of the algorithms proposed, we ran a set of experiments with the CIFAR-10 (Krizhevsky, 2009) and Tiny ImageNet datasets  (Le and Yang, 2015), using fully-connected and convolutional neural networks, specifically aiming to quantify the performance differences between the traditional (NP), layer-wise iterative (INP) and activity-based (ANP) formulations of NP as well as their decorrelated counterparts. These datasets were chosen as they have previously been shown to yield low performance when training networks using regular node perturbation, demonstrating its limitations (Hiratani et al., 2022). All experiments were repeated using three random seeds, after which performance statistics were aggregated. Further experimental details can be found in Appendix F.
3 Results
3.1 INP and ANP align better with the true gradient
Figure 1: Left: Angles between BP’s weight update and an update calculated by NP, ANP and INP as a function of the number of noise iterations (left) and the cost of each algorithm in forward passes (right). Note that the network is not updated across the ‘Noise iterations’ dimension, but instead multiple noise samples are propagated through a fixed network and their updates averaged. This is computed for a fully-connected, 3 hidden layer, network with leaky ReLU nonlinearities. Right: Number of forward passes for NP, ANP and INP.
When measuring the angles between the update vectors of various methods, we can observe that the INP method is by far the most well-aligned in its updates with respect to backpropagation, followed by ANP and closely thereafter by NP. These results align with the theory laid out in the methods section of this work. Appendix G shows the same data as a function of the number of noisy forward passes.
Note that for all of these algorithms, alignment with BP updates improves when updates are averaged over more samples of noise. Therefore, it is the ranking of the angles between the NP algorithms that is of interest here, not the absolute value of the angle itself, as all angles would improve with more noise samples. Appendix H shows that our results do not depend strongly on the noise variance $\sigma^{2}$, producing similar results when varying $\sigma^{2}$ by a few orders of magnitude.
3.2 Decorrelation improves convergence
Figure 2: Top panel: Performance of (D)NP and (D)BP on CIFAR-10 when training a single-layer architecture. Bottom panel: Performance of (D)NP, (D)ANP, (D)INP and (D)BP on CIFAR-10 in a three-hidden layer network. For both figures, curves report mean train and test accuracy. Shaded areas indicate the maximal and minimal accuracy obtained for three random seeds. Note that all NP methods have an equivalent formulation in a single-layer network.
To assess the impact of decorrelation on NP’s performance, we studied a single-layer network trained with NP, DNP, BP and DBP. Note that the different formulations of NP are identical in a single-layer networks where there is no impact on activities in the layer from previous layers. Figure 2 (Top panel) indicates that, in a single-layer context, DNP outperforms NP significantly and performs slightly better than BP, with DBP performing better still.
It appears that part of the benefit of decorrelation is due to a lack of correlation in the input unit features and a corresponding ease in credit assignment without confound.
We proceed by determining the performance of the different algorithms in multi-layer neural networks. See Appendix I, Table LABEL:tab:peakacc for peak accuracies.
Figure 2 (Bottom panel) shows the performance when training a multi-layer fully connected network. We see all decorrelated algorithms convincingly outperforming their non-decorrelated counterparts. We also see (D)INP outperform (D)ANP and (D)NP, almost keeping up with BP. Performance benefits of (D)ANP over (D)NP are not as clear in these results. Note that DNP performs much better in a three-hidden-layer network than in the single-layer networks, indicating that DNP does facilitate multi-layer credit assignment much more than regular NP, which actually performs worse in the deeper network. As an additional demonstration that our results generalize to other datasets, Appendix J reports results on the SARCOS dataset (Vijayakumar and Schaal, 2000). Appendix  K explores the scaling of DNP, DANP and DINP to networks with 6 and 9 hidden layers as well as a convolutional neural network trained on CIFAR-10.
3.4 Node perturbation allows learning in noisy systems
$\delta{\mathbf{a}_{l}}={\tilde{\mathbf{a}}}^{(1)}_{l}-\tilde{\mathbf{a}}^{(2)}_{l}$
Figure 3: Performance of DNP and DANP when using two noisy network passes. Results are reported for a three-layer fully-connected network trained for CIFAR-10 classification (comparable to Figure 2, bottom panel). Curves report mean train and test accuracy. Shaded areas indicate the maximal and minimal accuracy obtained for three random seeds.
3.5 Resampling allows for scaling to larger systems
To further investigate the capabilities of our approach, we apply BP, DNP, DANP and DINP to the more challenging Tiny ImageNet dataset Le and Yang (2015) using a deeper architecture consisting of three convolutional layers and three fully connected layers. The Tiny ImageNet dataset has 200 classes and thus has a larger output space than CIFAR-10. Larger output spaces are known to be challenging for traditional NP-style algorithms (Hiratani et al., 2022). RGB-values for the images, which are ImageNet images downsized to $64\times 64$ pixels, were standardized as a preprocessing step. For network architecture and learning rates, see Appendix F.
Figure 4 shows that, in their naive implementation, DNP, DANP and DINP do not appreciably learn this task. We demonstrate that, for such a difficult task, this is due to the accuracy of the gradient estimation. To show this, we resampled the random noise 10, 30 and 100 times before applying each update, effectively carrying out 10/30/100 noisy passes to compare against each single clean pass. For DNP, only an experiment with 100 iterations was included, as this algorithm did not match the performance of DINP and DANP. Results again demonstrate that both DANP and DINP improve meaningfully on DNP in deeper networks. The additional noise sampling comes at significant computational expense but could in principle be parallelized with sufficiently powerful hardware. The results indeed show that, as our theory suggests, sampling the noise distribution multiple times per update step significantly improves alignment with backpropagation, with DINP with 100 samples per update closely approximating BP’s learning trajectory. See Appendix I, Table LABEL:tab:peakacc_tin for peak accuracies.
4 Discussion
Though the presented methods for effective noise-based learning show a great deal of promise, there are a number of additional research steps to be taken. Testing the scalability of these approaches to tasks of greater complexity is important, as are their application to other network architectures such as residual networks, recurrent networks and attention-based models. Training accuracy, of even the best NP variants, also show some lag behind backpropagation’s accuracy achievements in deeper networks. Further investigation is needed to determine how to bridge this gap and if it has implications for harder tasks.
Our scaling results do demonstrate that by resampling the noise we achieve better and better alignment with the gradient direction, by virtue of the formal correspondence between directional derivatives and our formulations of node perturbation. In this sense, our approach can be viewed as an alternative to forward accumulation, where we exchange memory complexity (maintaining all partial derivatives) for time complexity (repeated forward passes with resampled noise vectors).
Appendix A Weights propagate noise correlations
Traditional NP updates have a number of sources of error. One reason is that use of the injected noise $\mathbf{\epsilon}_{l}$ to compute the learning signal ignores how the output activity of layer $l$ is not only impacted by the noise injected directly into it, but also by the cumulative effect of perturbations added to previous layers. If the noise is randomly sampled from a Gaussian distribution with mean zero, one might be tempted to simply assume that the perturbations from previous layers cancel out in expectation, but this assumption ignores the correlations introduced into these perturbations by the network’s weight matrices at all preceding layers. Multiplying a random vector by a non-orthogonal matrix will introduce a non-random covariance structure into it.
To see why, consider a transformation $\mathbf{W}\mathbf{x}=\mathbf{y}$, where $\mathbf{W}$ is a randomly initialized weight matrix and $\mathbf{x}$ is an uncorrelated input vector for which it holds that $\langle\mathbf{x}\mathbf{x}^{\top}\rangle=\mathbf{I}$ with $\mathbf{I}$ the identity matrix. The covariance matrix of $\mathbf{y}$ can be described as
$\textrm{cov}_{\mathbf{y}}$ $=\langle\mathbf{y}\mathbf{y}^{\top}\rangle=\langle(\mathbf{W}\mathbf{x})(\mathbf{W}\mathbf{x})^{T}\rangle=\langle(\mathbf{W}\mathbf{x})(\mathbf{x}^{\top}\mathbf{W}^{\top})\rangle$
$=\langle\mathbf{W}\mathbf{x}\mathbf{x}^{\top}\mathbf{W}^{\top}\rangle=\mathbf{W}\langle\mathbf{x}\mathbf{x}^{\top}\rangle\mathbf{W}^{\top}=\mathbf{W}\mathbf{I}\mathbf{W}^{\top}=\mathbf{W}\mathbf{W}^{\top}$
Therefore, any matrix for which $\mathbf{W}\mathbf{W}^{\top}\neq\mathbf{I}$ will add some covariance structure into output vector $\mathbf{y}$, even when input vector $\mathbf{x}$ is uncorrelated.
Appendix B The directional derivative as a measure of the gradient
$\mathbf{g}_{l}=N_{l}\left\langle{\nabla}_{\mathbf{v}}\mathcal{L}\frac{\mathbf{v}_{l}}{||\mathbf{v}||}\right\rangle_{\mathbf{v}}$
$\nabla_{\mathbf{v}}\mathcal{L}=\nabla\mathcal{L}^{\top}\frac{\mathbf{v}}{||\mathbf{v}||}\,.$
$\mathbf{g}$ $=N\left\langle{\nabla}_{\mathbf{v}}\mathcal{L}\frac{\mathbf{v}^{\top}}{||\mathbf{v}||}\right\rangle_{\mathbf{v}}^{\top}=N\left\langle\nabla\mathcal{L}^{\top}\frac{\mathbf{v}}{||\mathbf{v}||}\frac{\mathbf{v}^{\top}}{||\mathbf{v}||}\right\rangle_{\mathbf{v}}^{\top}=N\left\langle\frac{\mathbf{v}\mathbf{v}^{\top}}{||\mathbf{v}||^{2}}\right\rangle_{\mathbf{v}}^{\top}\nabla\mathcal{L}\,.$
$\mathbf{g}$ $=N\left\langle\frac{\mathbf{v}\mathbf{v}^{\top}}{||\mathbf{v}||^{2}}\right\rangle_{\mathbf{v}}^{\top}\nabla\mathcal{L}\propto N\frac{\mathbf{I}}{N}\nabla\mathcal{L}=\nabla\mathcal{L}\,.$
Note that normalization of every vector means that the correlations $\left\langle\ifrac{\mathbf{v}\mathbf{v}^{\top}}{||\mathbf{v}||^{2}}\right\rangle$ are not changed in sign but only in scale, as the noise vectors are now all scaled to lie on a unit sphere (without any rotation). In our specific case of a diagonal correlation matrix, this expectation is also therefore diagonal and all diagonal elements are equal. The average value of these diagonal elements is then simply the variance ${1}/{N}$ of a random unit vector in $N$ dimensional space. For other noise distributions, in which there does exist cross-correlation between the noise elements, this is no longer an appropriate treatment. Instead, one would have to multiply by the inverse of the correlation matrix to fully recover the gradient values.
Appendix C Derivation of activity-based node perturbation
$\Delta\mathbf{W}_{l}^{\textrm{INP}}=N_{l}\>\delta\mathcal{L}\frac{\mathbf{v}_{l}}{||\mathbf{v}||^{2}}\>\mathbf{x}_{l-1}^{\top}\,.$
This rule has been derived in the main text in order to optimally make use of noise to determine the directional derivative with respect to a single layer of a deep network, $l$, while all other layers receive no noise. For this purpose, the noise vector is defined such that $\mathbf{v}_{k}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I}_{k})$ if $k=l$ and $\mathbf{v}_{k}=\mathbf{0}$ otherwise. In moving from INP to ANP, we aim to fulfill a number of conditions:
  1. 1.
    Update a whole network in one pass rather than per layer.
  2. 2.
    Allow updating without explicit access to the noise at every layer, but instead access only to the activity difference.
$\Delta\mathbf{W}_{l}^{\textrm{ANP}}=N\>\delta{\mathcal{L}}\frac{\delta{\mathbf{a}_{l}}}{||\delta\mathbf{{a}}||^{2}}\>\mathbf{x}_{l-1}^{\top}$
which can be computed via two forward passes only.
Appendix D Gradient Alignment of NP vs ANP
In this section, we analyze the deviation of NP’s and ANP’s updates from the true gradient in a simple two-layer neural network, focusing on the effects of noise propagation through the network’s layers.
In practice, perturbations applied to networks cannot be infinitely small and are even limited by precision of floating point representations to be larger than some minimum size. Nonetheless, we here assume that perturbations are sufficiently small that networks can be assumed to have been linearized.
Consider a two-layer neural network with weight matrices $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$, no biases, and a nonlinear activation function $f(\cdot)$ applied after the first layer. The network’s output, given an input $x$, is expressed as:
$\mathbf{y}_{2}=\mathbf{W}_{2}f(\mathbf{y}_{1})=\mathbf{W}_{2}f(\mathbf{W}_{1}\mathbf{x})\,.$
Noise is introduced into the network as small perturbations $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ at the outputs of the first and second layers, respectively. With this perturbation, the perturbed output becomes:
$\tilde{\mathbf{y}}_{2}=\mathbf{W}_{2}\big{(}f(\mathbf{y}_{1}+\mathbf{v}_{1})\big{)}+\mathbf{v}_{2}=\mathbf{W}_{2}\big{(}f(\mathbf{W}_{1}\mathbf{x}+\mathbf{v}_{1})\big{)}+\mathbf{v}_{2}\,.$
Given an arbitrary loss function $L$, the loss differential is given by $L(\tilde{\mathbf{y}}_{2})-L(\mathbf{y}_{2})$. NP’s updates are based on a product between this loss differential and the noise. Examining this for matrix $\mathbf{W}_{2}$, this corresponds to a product between the noise differential and vector $\mathbf{v}_{2}$. To fully show the impact of this update, we expand the loss differential by a first-order Taylor expansion.
As such,
$\langle\Delta\mathbf{W}_{2}^{\text{NP}}\rangle$ $=\frac{1}{\sigma^{2}}\left\langle\delta L\cdot\mathbf{v}_{2}\right\rangle f(\mathbf{y}_{1})^{\top}$
$=\frac{1}{\sigma^{2}}\left\langle\mathbf{v}_{2}\,(L(\tilde{\mathbf{y}}_{2})-L(\mathbf{y}_{2}))^{\top}\right\rangle f(\mathbf{y}_{1})^{\top}$
$\approx\frac{1}{\sigma^{2}}\left\langle\mathbf{v}_{2}\,(\nabla_{\mathbf{y}_{2}}L^{\top}(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2}))^{\top}\right\rangle f(\mathbf{y}_{1})^{\top}$
$\approx\frac{1}{\sigma^{2}}\left\langle\mathbf{v}_{2}\,(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})^{\top}\nabla_{\mathbf{y}_{2}}L\right\rangle f(\mathbf{y}_{1})^{\top}$
$\approx\frac{1}{\sigma^{2}}\left\langle\mathbf{v}_{2}\,(\mathbf{v}_{2}+\nabla_{\mathbf{y_{1}}}\mathbf{y_{2}}\circ\mathbf{v}_{1})^{\top}\right\rangle\nabla_{\mathbf{y}_{2}}L\,f(\mathbf{y}_{1})^{\top}$
$\approx\frac{1}{\sigma^{2}}\left\langle\mathbf{v}_{2}\mathbf{v}_{2}^{\top}+\mathbf{v}_{2}\mathbf{v}_{1}^{\top}\circ\nabla_{\mathbf{y_{1}}}\mathbf{y_{2}}^{\top}\right\rangle\nabla_{\mathbf{y}_{2}}L\,f(\mathbf{y}_{1})^{\top}$
In regular node perturbation, it is thereafter supposed that the terms in the angular brackets cancel down to the identity, $\langle\mathbf{v}_{2}\mathbf{v}_{2}^{\top}\rangle=\mathbf{I}$ and $\langle\mathbf{v}_{2}\mathbf{v}_{1}^{\top}\rangle=\mathbf{0}$, given our initial assumptions about noise sampling. However, in the case in which only a single (or relatively few) noise samples are used to compute an NP update, both of the terms related to the correlation between nodes noise at the second layer, $\langle\mathbf{v}_{2}\mathbf{v}_{2}^{\top}\rangle$, and the correlation between node noise between the first and second layers, $\langle\mathbf{v}_{2}\mathbf{v}_{1}^{\top}\rangle$, contribute an error to this update. Most notably, any error in the first term contributes to the gradient descent update being multiplied by a positive semi-definite matrix, which still ensures that the loss will be reduced. However any error in the second term, i.e. $\langle\mathbf{v}_{2}\mathbf{v}_{1}^{\top}\rangle\neq\mathbf{0}$, is not positive semi-definite and can throw off the update of NP away from reducing the loss.
Carrying out the same analysis for the ANP rule,
$\langle\Delta\mathbf{W}_{2}^{\text{ANP}}\rangle$ $=\Big{\langle}\gamma\big{(}L(\tilde{\mathbf{y}}_{2})-L(\mathbf{y})\big{)}\cdot(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})\Big{\rangle}f(\mathbf{y}_{1})^{\top}$
$\approx\Big{\langle}(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})\Big{[}\nabla_{\mathbf{y}_{2}}L^{\top}(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})\Big{]}^{\top}\Big{\rangle}f(\mathbf{y}_{1})^{\top}$
$\approx\Big{\langle}(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})^{\top}\nabla_{\mathbf{y}_{2}}L\Big{\rangle}f(\mathbf{y}_{1})^{\top}$
$\approx\Big{\langle}(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})^{\top}\Big{\rangle}\nabla_{\mathbf{y}_{2}}Lf(\mathbf{y}_{1})^{\top}$
where $\gamma$ represents the normalizing factor required for the update rule. ANP’s relationship to the true gradient is far simpler than that of NP, only depends upon the output activations difference, and reduces the algorithms bias down to the correlation in overall activity difference at this layer ($\langle(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})^{\top}\rangle$). This correlation will strictly contribute a positive semi-definite matrix multiplication against the gradient update (regardless of how many noise samples are used for the averaging), thus ensuring that the loss is nonetheless reduced. It is also a much more interpretable bias term which could even be targeted directly to be made diagonal if a practitioner was interested in sampling to ensure this property.
Note that the comparisons in this section lean upon on the well known property that positive semi-definite matrices never modify a vector to be more than $90^{\circ}$ away from it’s initial orientation.
Appendix E Algorithm pseudocode
@@
         $\delta\mathcal{L}=(\mathbf{t}-\mathbf{\tilde{x}}_{L})^{2}-(\mathbf{t}-\mathbf{x}_{L})^{2}$triangleright Compute loss difference
@@
             $\mathbf{W}_{l}\leftarrow\mathbf{W}_{l}-\eta N\,\delta\mathcal{L}\,\frac{\mathbf{\tilde{a}}_{l}-\mathbf{a}_{l}}{||\delta\mathbf{{a}}||^{2}}\left({\mathbf{x}^{\star}_{l-1}}\right)^{\top}$triangleright Update weight matrix
             $\mathbf{R}_{l}\leftarrow\mathbf{R}_{l}-\alpha\left(\mathbf{x}^{\star}_{l}\left(\mathbf{x}^{\star}_{l}\right)^{\top}-\textrm{diag}\left(\left(\mathbf{x}^{\star}_{l}\right)^{2}\right)\right)\mathbf{R}_{l}$triangleright Update decorrelation matrix
@@
Appendix F Experimental details

### v7

@@ 2.2 Increasing perturbation learning efficiency through decorrelation
Decorrelation effectively aligns the gradient updates more closely with those of the natural gradient (Ahmad, 2024) and can be combined with any of the formulations described in the previous sections by replacing $\mathbf{x}_{l}$ in Eq. 1 with $\mathbf{x}_{l}^{*}$. We will use DBP, DNP, DINP and DANP when referring to the decorrelated versions of the described learning rules.
Algorithm 1, describes the decorrelated activity-based node perturbation procedure, which we will argue is of particular interest for implementations on unconventional computing hardware.
@@
    $\delta\mathcal{L}=(\mathbf{t}-\mathbf{\tilde{x}}_{L})^{2}-(\mathbf{t}-\mathbf{x}_{L})^{2}$ $\triangleright$ Compute loss difference
@@
[…]
     $\mathbf{R}_{l}\leftarrow\mathbf{R}_{l}-\alpha\left(\mathbf{x}^{\star}_{l}\left(\mathbf{x}^{\star}_{l}\right)^{\top}-\textrm{diag}\left(\left(\mathbf{x}^{\star}_{l}\right)^{2}\right)\right)\mathbf{R}_{l}$ $\triangleright$ Update decorrelation matrix
@@
2.3 Experimental validation
Our experiment in Section 3.1 uses a synthetic input–output mapping, where input and target vectors $\mathbf{x}$ are sampled from a Gaussian distribution $\mathcal{N}(\mathbf{0},\mathbf{1})$. When computing updates, perturbations are applied by adding small random noise to layer activations, where each perturbation vector is drawn from $\mathcal{N}(\mathbf{0},\sigma^{2})$ with $\sigma^{2}=10^{-6}$. These updates are then compared to updates computed by BP to determine their correspondence by measuring the angle, measured as $\cos\theta=\frac{\mathbf{u}\cdot\mathbf{v}}{|\mathbf{u}||\mathbf{v}|}$, where $\theta$ is the angle, $\mathbf{u}$ and $\mathbf{v}$ are the vectors, and $|\cdot|$ denotes the Euclidean norm. For all other experiments in this work using NP, ANP or INP $\sigma^{2}=10^{-6}$ was also used when sampling the perturbations. See Appendix C for more details on how this value was determined.
To measure the performance of the algorithms proposed, we ran a set of experiments with the CIFAR-10 (Krizhevsky, 2009) and Tiny ImageNet datasets (Le and Yang, 2015), using fully-connected and convolutional neural networks, specifically aiming to quantify the performance differences between the traditional (NP), layer-wise iterative (INP) and activity-based (ANP) formulations of NP as well as their decorrelated counterparts. These datasets were chosen as they have previously been shown to yield low performance when training networks using regular node perturbation, demonstrating its limitations (Hiratani et al., 2022). All experiments were repeated using three random seeds, after which performance statistics were aggregated. Further experimental details can be found in Appendix D. Additionally, we performed experiments on the SARCOS dataset (Vijayakumar and Schaal, 2000), demonstrating the application of our work in a robotics setting.
3 Results
3.1 INP and ANP align better with the true gradient
Figure 1: Left: Angles between BP’s weight update and an update calculated by NP, ANP and INP as a function of the number of noise iterations. Note that the network is not updated across the ‘Noise iterations’ dimension, but instead multiple noise samples are propagated through a fixed network and their updates averaged. This is computed for a fully-connected, 3 hidden layer, network with leaky ReLU nonlinearities. Right: Number of forward passes for NP, ANP and INP.
When measuring the angles between the update vectors of various methods, we can observe that the INP method is by far the most well-aligned in its updates with respect to backpropagation, followed by ANP and closely thereafter by NP. These results align with the theory laid out in the methods section of this work. Appendix E shows the same data as a function of the number of noisy forward passes.
Note that for all of these algorithms, alignment with BP updates improves when updates are averaged over more samples of noise. Therefore, it is the ranking of the angles between the NP algorithms that is of interest here, not the absolute value of the angle itself, as all angles would improve with more noise samples. Appendix C shows that our results do not depend strongly on the noise variance $\sigma^{2}$, producing similar results when varying $\sigma^{2}$ by a few orders of magnitude.
3.2 Decorrelation improves convergence
To assess the impact of decorrelation on NP’s performance, we studied a single-layer network trained with NP, DNP, BP and DBP. Note that the different formulations of NP are identical in a single-layer networks where there is no impact on activities in the layer from previous layers.
Figure 2: Performance of (D)NP and (D)BP on CIFAR-10 when training a single-layer architecture.
Figure 2 indicates that, in a single-layer context, DNP outperforms NP significantly and performs slightly better than BP, with DBP performing better still. It appears that part of the benefit of decorrelation is due to a lack of correlation in the input unit features and a corresponding ease in credit assignment without confound.
Figure 3: Performance of (D)NP, (D)ANP, (D)INP and (D)BP on CIFAR-10 in a three-hidden layer network. For both figures, curves report mean train and test accuracy. Shaded areas indicate the maximal and minimal accuracy obtained for three random seeds. Note that all NP methods have an equivalent formulation in a single-layer network.
We proceed by determining the performance of the different algorithms in multi-layer neural networks. See Appendix F, Table 3 for peak accuracies. Figure 3 shows the performance when training a multi-layer fully connected network. We see all decorrelated algorithms convincingly outperforming their non-decorrelated counterparts. We also see (D)INP outperform (D)ANP and (D)NP, almost keeping up with BP. Performance benefits of (D)ANP over (D)NP are not as clear in these results. Note that DNP performs much better in a three-hidden-layer network than in the single-layer networks, indicating that DNP does facilitate multi-layer credit assignment much more than regular NP, which actually performs worse in the deeper network.
3.4 Resampling allows for scaling to larger systems
To further investigate the capabilities of our approach, we apply BP, DNP, DANP and DINP to the more challenging Tiny ImageNet dataset (Le and Yang, 2015) using a deeper architecture consisting of three convolutional layers and three fully connected layers. The Tiny ImageNet dataset has 200 classes and thus has a larger output space than CIFAR-10. Larger output spaces are known to be challenging for traditional NP-style algorithms (Hiratani et al., 2022). RGB-values for the images, which are ImageNet images downsized to $64\times 64$ pixels, were standardized as a preprocessing step. For network architecture and learning rates, see Appendix D.
Figure 4 shows that, in their naive implementation, DNP, DANP and DINP do not appreciably learn this task. We demonstrate that, for such a difficult task, this is due to the accuracy of the gradient estimation. To show this, we resampled the random noise 10, 30 and 100 times before applying each update, effectively carrying out 10/30/100 noisy passes to compare against each single clean pass. For DNP, only an experiment with 100 iterations was included, as this algorithm did not match the performance of DINP and DANP. Results again demonstrate that both DANP and DINP improve meaningfully on DNP in deeper networks. The additional noise sampling comes at significant computational expense but could in principle be parallelized with sufficiently powerful hardware. The results indeed show that, as our theory suggests, sampling the noise distribution multiple times per update step significantly improves alignment with backpropagation, with DINP with 100 samples per update closely approximating BP’s learning trajectory. See Appendix F, Table LABEL:tab:peakacc_tin for peak accuracies.
3.5 Robot control task
As an additional demonstration that our results generalize to other datasets with applications in robotics, we report results on the SARCOS dataset (Vijayakumar and Schaal, 2000).
Figure 5 shows train and test loss for (D)NP, (D)ANP and (D)BP for the SARCOS dataset (Vijayakumar and Schaal, 2000). The task is to solve inverse dynamics problem for a seven degrees-of-freedom SARCOS anthropomorphic robot arm by mapping from a 21-dimensional input space (7 joint positions, 7 joint velocities, 7 joint accelerations) to the corresponding 7 joint torques. The network is a fully connected architecture with two 32-unit hidden layers and was trained with MSE loss and the leaky ReLU activation function. The network was trained for 500 epochs and results were averaged over three random seeds. A learning rate search for each algorithm was started at $10^{-7}$, doubling the learning rate until the algorithm became unstable. A decorrelation learning rate of $10^{-4}$ was used for DNP and DANP and $10^{-3}$ for DBP.
3.6 Node perturbation allows learning in noisy systems
$\delta{\mathbf{a}_{l}}={\tilde{\mathbf{a}}}^{(1)}_{l}-\tilde{\mathbf{a}}^{(2)}_{l}$
Figure 6: Performance of DNP and DANP when using two noisy network passes. Results are reported for a three-layer fully-connected network trained for CIFAR-10 classification (comparable to Figure 3). Curves report mean train and test accuracy. Shaded areas indicate the maximal and minimal accuracy obtained for three random seeds.
4 Discussion
It is important to note that adding decorrelation matrices to a neural network does increase the amount of computation required in the forward pass. Appendix G provides details on the exact amount of additional computation and memory required when running BP, DBP or DANP. It should be noted, however, that several computational tricks are available that will significantly reduce said computational burden. These methods have not been applied in the current work, due to the small architectures trained, but would likely benefit runtimes in larger networks. First, the decorrelation matrix $\mathbf{R}$ can be fused with forward weights $\mathbf{W}$, such that $\mathbf{A}=\mathbf{W}\mathbf{R}$ and $\mathbf{x}=\mathbf{A}\mathbf{z}$, moving the multiplication by $\mathbf{R}$ to $\mathbf{W}$, which does not contain a batch dimension, significantly reducing computation in the forward pass. Second, the update of $\mathbf{R}$ need not be estimated using the entire mini-batch, but can also be estimated using a heavily downsampled mini-batch (Dalm et al., 2024). Third, in the case of convolutional layers, only the local image patch on which the convolutional kernel operates could be decorrelated (Dalm et al., 2024). As the only correlations that matter when updating a convolutional kernel are the local correlations within its receptive field, this should not diminish performance gains, while greatly reducing computational overhead.
Testing the scalability of these approaches to tasks of greater complexity is also important, as are their application to other network architectures such as residual networks, recurrent networks and attention-based models. Training accuracy, of even the best NP variants, also show some lag behind backpropagation’s accuracy achievements in deeper networks. Further investigation is needed to determine how to bridge this gap and if it has implications for harder tasks. Our scaling results do demonstrate that by resampling the noise we achieve better and better alignment with the gradient direction, by virtue of the formal correspondence between directional derivatives and our formulations of node perturbation. In this sense, our approach can be viewed as an alternative to forward accumulation, where we exchange memory complexity (maintaining all partial derivatives) for time complexity (repeated forward passes with resampled noise vectors).
Acknowledgments and Disclosure of Funding
This publication is part of the project Dutch Brain Interface Initiative (DBI2) with project number 024.005.022 of the research programme Gravitation which is (partly) financed by the Dutch Research Council (NWO).
Appendix A The directional derivative as a measure of the gradient
[…]
$𝐯.\nabla_{\mathbf{v}}\mathcal{L}$$\nabla\mathcal{L}^{\top}\frac$$\mathbf$$|\mathbf{v}$||}\,.
[…]
Appendix B Derivation of activity-based node perturbation
[…]
This rule has been derived in the main text in order to optimally make use of noise to determine the directional derivative with respect to a single layer of a deep network, $l$, while all other layers receive no noise. For this purpose, the noise vector is defined such that $\mathbf{v}_{k}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I}_{k})$ if $k=l$ and $\mathbf{v}_{k}=\mathbf{0}$ otherwise. In moving from INP to ANP, we aim to fulfill the following conditions. First, update a whole network in one pass rather than per layer. Second, allow updating without explicit access to the noise at every layer, but instead access only to the activity difference.
[…]
which can be computed via two forward passes only.
Appendix C Impact of perturbation scale on performance
Appendix D Experimental details